OpenStudy (anonymous):

what percent of 1440 is 6

OpenStudy (anonymous):

6/1440 THEN DECIMAL INTO PERCENT

OpenStudy (anonymous):

\[\frac{ x }{ 100 } \times 1440 = 6\]

OpenStudy (anonymous):

then find x...

OpenStudy (anonymous):

x=6/1440*100

OpenStudy (candy13106):

ok so did u ever get the answer

OpenStudy (candy13106):

ok well i guess ur off

OpenStudy (anonymous):

Let x percent of 144O is 6. Now , x/100* 1440=6 => x=4.167

OpenStudy (candy13106):

no all wrong u lerarened it the wrong way wait ill tell u

OpenStudy (candy13106):

ok do.................

OpenStudy (candy13106):